The Baldwin's balance sheet has $88,312,000 in equity. Further, the company is expecting $3,000,000 in net income next year. Assuming no dividends are paid and no stock is issued, what would their Book Value be next year?
Select: 1
$53.27
$22.71
$23.48

$8.48

Company Andrews Baldwin Chester Digby Close $195,46 $19.28 $107.32 $30.28 Change $90.79 3,75 $32.21 $10.17 Shares MarketCap (SM) Book Value 1,893,587 $370 $81.02 $75 3,889,135 22.71 $39.58 1,824,138 $196 3,034,114 $92. $22.00 EPS $29.08 $0.12 $13.11 $2.00 Dividend $0.00 $0.00 $10.31 $2.04 Yield 0.0% 0.0% 9,6% 6.7% PIE 6,7 161,7 5,1
Answer #1

Ans : $ 23.48

Solution : Book Value Per Share = (ShareHolder Equity+ Net Income)/ No of Shares

For Current Year : Shareholder Equity = 88312000

                           Book Value Per Share = 22.71

                           No of shares = 88312000/22.71 = 3888683 shares (Approx)

For Next Year :

    Book Value Per Share = ( 88312000 + 3000000)/ 3888683 shares

                                      = 23.48 per share